1. If I have 6 moles of a gas at a pressure of 5.6 atm and a volume of 12 liters, what is the temperature?
2. If I have an unknown quantity of gas at a pressure of 1.8 atm, a volume of 41 liters, and a temperature of 87°C, how many moles of gas do I have?
3. If I have 20 moles of gas at a temperature of 67°C and a volume of 91.3 liters, what is the pressure of the gas?
4. If I place 3 moles of a gas in a container with a volume of 75 liters and at a temperature of 350 K, what is the pressure inside the container?
5. If I have 6.7 moles of gas at a pressure of 1.24 atm and at a temperature of 62°C, what is the volume of the container that the gas is in?
1
Expert's answer
2020-04-10T13:36:33-0400

1) PV=nRT

n=6mol

P=5,6atm

V=12l

R=0.0821

T=?

T=PV/nR= (5.6*12)/(6*0.0821)=67.2/0.4926=136K


Answer: T=136K


2) PV=nRT

P=1.8atm

V=41l

T=87°=357K

R=0.0821

n=?

n=PV/RT=(1.8*41)/(0.0821*357)=73.8/29.3=2.52mol


Answer: n=2.52mol


3) PV=nRT

n=20mol

T=67°=337K

V=91.3l

R=0.0821

P=?

P=nRT/V=(20*0.0821*337)/91.3=553.3/91.3=6.06atm


Answer: P=6.06atm


4) PV=nRT

n=3mol

V=75l

T=350K

R=0.0821

P=?

P=nRT/V=(3*0.0821*350)/75=86.2/75=1.149atm


Answer: P=1.149atm


5) PV=nRT

n=6.7mol

P=1.24atm

T=62°=332K

R=0.0821

V=?

V=nRT/P=(6.7*0.0821*332)/1.24=182.6/1.24=147l


Answer: V=147l